OpenStudy (anonymous):

How about this : (6xsquare-5x-bracket3x-(4xsquare-9) close bracket

OpenStudy (anonymous):

\[Could some one help me with this problem please show me how you worked \it out (6xsquare-5x \left[ \left[ 3x-(4xsquare-9(\left[ ? \right] \ ?) & ? \ ? & ?) \right] \right]\]